OpenStudy (anonymous):

simplify the rational expression q^2+11q+24/q^2-5q-24

jigglypuff314 (jigglypuff314):

factor the numerator and the denominator q^2+11q+24 = (q+3)(q+8) q^2-5q-24 = (q+3)(q-8)

OpenStudy (anonymous):

Factor the numerator and denominator of your fraction, then cancel like terms.